OpenStudy (lgbasallote):

A 0.4481-g sample of a compound known to contain only carbon, hydrogen, and oxygen was burned in oxygen to yield 0.8326 g of CO2 and 0.1705 g of H2O. What is the empirical formula of the compound? A) CHO B) C2H2O C) C3H3O2 D) C6H3O2 E) C3H6O2 i just need a GENERAL IDEA on what to do

OpenStudy (lgbasallote):

@sami-21 help

OpenStudy (anonymous):

Last time i saw my chemistry book was 4 years ago :P

OpenStudy (lgbasallote):

of what?

OpenStudy (anonymous):

is there any easy way than this one? http://answers.yahoo.com/question/index?qid=20080927180546AAWEvyP

OpenStudy (lgbasallote):

would be if that would be explained...

OpenStudy (unklerhaukus):

Find the stoichiometry of \[\text{C}_x\text H _y \text O_z+\text O_2\rightarrow\text {CO}_2+\text H_2\text O\]

OpenStudy (lgbasallote):

by that you mean?

OpenStudy (unklerhaukus):

determine the number of moles of the products

OpenStudy (lgbasallote):

so i have to balance first @UnkleRhaukus ?

OpenStudy (lgbasallote):

i have no idea what you meant @UnkleRhaukus ....honestly....

OpenStudy (unklerhaukus):

0.8326 g of CO_2 how many moles is this?

OpenStudy (lgbasallote):

i got 0.07

OpenStudy (lgbasallote):

oh wait...

OpenStudy (lgbasallote):

0.019

OpenStudy (unklerhaukus):

ok good, 0.02 moles of CO_2 how many moles of H_2O 0.1705 g

OpenStudy (lgbasallote):

weird... i got 0.0095

OpenStudy (unklerhaukus):

ok so we have \[\frac{0.448}{x\times12.0107+y\times1.0079+z\times15.9994}[\text C_x\text H_y \text O_z ]\]\[\rightarrow\]\[\frac{ 0.8326}{12.0107+2\times15.9994}[\text {CO}_2]+\frac{0.1705}{2\times1.0079+15.9994}[\text H_2\text O]\] solve for x,y,z

OpenStudy (unklerhaukus):

whops 0.448 should be 0.4481

OpenStudy (lgbasallote):

uhmm explain?

OpenStudy (lgbasallote):

oh nevermind

OpenStudy (unklerhaukus):

i went backwards one step because the accuracy looks like it is important

OpenStudy (lgbasallote):

so what happens next? cross multiplication?

OpenStudy (unklerhaukus):

you could do that , im gonna try simply those fractions again to high accuracy

OpenStudy (unklerhaukus):

\[\small \frac{0.448}{x\times12.0107+y\times1.0079+z\times15.9994}[\text C_x\text H_y \text O_z ]\rightarrow 0.01892[\text {CO}_2]+ 0.009464 [\text H_2\text O]\]

OpenStudy (unklerhaukus):

i forgot how imprecise theses calculations are comparing 0.01892 & 0.009464 we see the moles of CO_2 is approximately twice the moles of H_2O

OpenStudy (lgbasallote):

i think this is starting to get more complicated than it's supposed to be o.O

OpenStudy (unklerhaukus):

multiply by one hundred

OpenStudy (unklerhaukus):

\[ \frac{44.8}{x\times12.0107+y\times1.0079+z\times15.9994}[\text C_x\text H_y \text O_z ]\rightarrow 2[\text {CO}_2]+ 1 [\text H_2\text O]\]

OpenStudy (unklerhaukus):

determine the stoichiometry by balancing the atomic species, for some reason i left oxygen off the LHS,

OpenStudy (unklerhaukus):

\[\text{C}_x\text H _y \text O_z+\text O_2\rightarrow2\text {CO}_2+1\text H_2\text O\] balance the carbons first

OpenStudy (unklerhaukus):

on the RHS \[2\text C : 2\text H : 5\text O\]

OpenStudy (unklerhaukus):

\[2\text C_x\text H_x \text O_{z}+\text O_2\rightarrow2\text {CO}_2+1\text H_2\text O\]\[4\text C_x\text H_x \text O_{z}+\text O_2\rightarrow4\text {CO}_2+2\text H_2\text O\]

OpenStudy (unklerhaukus):

\[4\text C\text H \text O+\text O_2\rightarrow4\text {CO}_2+2\text H_2\text O\]